# A general framework of Riemannian adaptive optimization methods with a convergence analysis

### v1

@@ 4 Numerical Experiments
We experimented with both constant and diminishing step sizes. For each algorithm, we searched in the set $\{10^{-1},10^{-2},\ldots,10^{-8}\}$ for the best initial step size $\alpha$ ( both constant and diminishing). Note that the constant (resp. diminishing) step size was determined to be $\alpha_{k}=\alpha$ (resp. $\alpha_{k}=\alpha/\sqrt{k}$) for all $k\geq 1$. The experiments used a MacBook Air (M1, 2020) and the macOS Monterey version 12.2 operating system. The algorithms were written in Python 3.12.1 with the NumPy 1.26.0 package and the Matplotlib 3.9.1 package. The Python implementations of the methods used in the numerical experiments are available at https://github.com/iiduka-researches/202408-adaptive.
@@ 4.1 Principal component analysis
on the Stiefel manifold $\operatorname*{St}(p,n)$. Therefore, the PCA problem can be considered to be optimization problem on the Stiefel manifold.
In the experiments, we set $p$ to $10$ and the batch size $b$ to $2^{10}$. We used the QR-based retraction on the Stiefel manifold $\operatorname*{St}(p,n)$ (Absil et al., 2008, Example 4.1.3), which is defined by
@@
We evaluated the algorithms on training images of the MNIST dataset (LeCun et al., 1998) and the COIL100 dataset (Nene et al., 1996). The MNIST dataset contains 60,000 $28\times 28$ gray-scale images of handwritten digits. We transformed every image into a 784-dimensional vector and normalized its pixel values to lie in the range of $[0,1]$. Thus, we set $N=60000$ and $n=784$. The COIL100 dataset contains 7,200 normalized color camera images of the 100 objects taken from different angles. As in the previous study (Kasai et al., 2019), we resized them to 32$\times$32 pixels. Thus, we set $N=7200$ and $n=1024$.
Figure 1(a) (resp. Figure 1(b)) shows the performances of the algorithms with a constant (resp. diminishing) step size for the objective function values defined by (8) with respect to the number of iterations on the MNIST dataset, while Figure 3(a) and 3(b) present those on the COIL100 dataset. Figure 2(a) (resp. 2(b)) presents the performances of the algorithms with a constant (resp. diminishing) step size for the norm of the gradient of objective function defined by (8) with respect to the number of iterations on the MNIST dataset, while Figure 4(a) and 4(b) present those on the COIL100 dataset. The experiments were performed for three random initial points, and the thick line plots the average of all experiments. The area bounded by the maximum and minimum values is painted the same color as the corresponding line.
@@
Figure 1: Objective function value defined by (8) versus number of iterations on the MNIST datasets.
@@
Figure 2: Norm of the gradient of objective function defined by (8) versus number of iterations on the MNIST datasets.
@@
Figure 3: Objective function value defined by (8) versus number of iterations on the COIL100 datasets.
@@
Figure 4: Norm of the gradient of objective function defined by (8) versus number of iterations on the COIL100 datasets.
Figure 1(a) indicates that RAdam and RAMSGrad (Algorithm 2) performed comparably to RASA-LR in the sense of minimizing the objective function value. Figure 1(b) indicates that RAdam and RAMSGrad (Algorithm 2) outperformed RASA-L and RASA-R. Figure 2(a) shows that RAMSGrad (Algorithm 2) performed better than RASA-LR in the sense of minimizing the full gradient norm of the objective function. Figure 3(a) indicates that RAdam and RAMSGrad (Algorithm 2) had the best performance in the sense of minimizing the objective function value. Figure 3(b) indicates that RAdam and RAMSGrad (Algorithm 2) performed comparably to RASA-LR. Figure 4(a) shows that RAdam had the best performance in the sense of minimizing the full gradient norm of the objective function. Figure 4(b) indicates that RAdam and RAMSGrad (Algorithm 2) performed comparably to RASA-R and RASA-LR.
@@ 4.2 Low-rank matrix completion
We evaluated the algorithms on the MovieLens-1M [^1] https://grouplens.org/datasets/movielens/ datasets (Harper & Konstan, 2015) and the Jester [^2] https://grouplens.org/datasets/jester datasets for recommender systems. The MovieLens-1M datasets contains 1,000,209 ratings given by 6,040 users on 3,952 movies. Thus, we set $N=3952$ and $n=6040$. The Jester datasets contains ratings of 100 jokes given by 24,983 users with scores from $-10$ to 10. Thus, we set $N=24983$ and $n=100$.
@@
Figure 5(a) (resp. Figure 5(b)) shows the performances of the algorithms with a constant (resp. diminishing) step size for objective function values defined by (10) with respect to the number of iterations on the MovieLens-1M dataset, while Figure 7(a) and 7(b) present those on the Jester dataset. Figure 6(a) (resp. 6(b)) shows the performances of the algorithms with a constant (resp. diminishing) step size for the norm of the gradient of the objective function defined by (8) with respect to the number of iterations on the MovieLens-1M dataset, while Figure 7(a) and 7(b) present those on the Jester dataset. The experiments were performed for three random initial points, and the thick line plots the average results of all experiments. The area bounded by the maximum and minimum values is painted the same color as the corresponding line.
@@
Figure 5: Objective function value defined by (10) versus number of epochs (iterations for the entire dataset) on the MovieLens-1M datasets.
@@
Figure 6: Norm of the gradient of objective function defined by (10) versus number of epochs (iterations for the entire dataset) on the MovieLens-1M datasets.
@@
Figure 7: Objective function value defined by (10) versus number of epochs (iterations for the entire dataset) on the Jester datasets.
@@
Figure 8: Norm of the gradient of objective function defined by (10) versus number of epochs (iterations for the entire dataset) on the Jester datasets.
Figure 5(a) indicates that RAMSGrad (Algorithm 2) performed better than RASA-L and RASA-LR in the sense of minimizing the objective function value. Figure 5(b) shows that RAdam and RAMGRad (Algorithm 2) performed comparably to RASA-L and RASA-LR. Figure 6(a) indicates that RAdam performed comparably to RASA-R in the sense of minimizing the full gradient norm of the objective function. Figure 6(b) shows that RAdam outperformed RASA-R. Figure 7(a) and 7(b) indicate that RAMSGrad (Algorithm 2) performed better than RASA-LR in the sense of minimizing the objective function value. Figure 8(a) indicates that RAdam performed comparably to RASA-L and RASA-R in the sense of minimizing the full gradient norm of the objective function. Figure 8(b) shows that RAMSGrad (Algorithm 2) outperformed RASA-L and RASA-R.
@@ 5 Conclusion
This paper proposed a general framework of Riemannian adaptive optimization methods, which encapsulates several stochastic optimization algorithms on Riemannian manifolds. The framework incorporates the mini-batch strategy often used in deep learning. We also proposed AMSGrad that works on embedded submanifolds of Euclidean space within our framework. In addition, we gave convergence analyses that are valid for both a constant and diminishing step size. The analyses also revealed the relationship between the convergence rate and mini-batch size. We numerically compared the AMSGrad with the existing algorithms by applying them to the principal component analysis and the low-rank matrix completion problem, which can be considered to be Riemannian optimization problems. Numerical experiments showed that the proposed method performs well against PCA. RAdam and RAMSGrad performed well for constant and diminishing step sizes especially on the COIL100 dataset.
Appendix A Proof of Lemma 3.2
@@
$\mathbb{E}_{k}\left[\operatorname*{grad}f_{B_{k}}(x_{k})\right]=\frac{1}{b}\sum_{i=1}^{b}\mathbb{E}_{k}\left[\operatorname*{grad}f_{s_{k,i}}(x_{k})\right]=\operatorname*{grad}f(x_{k}).$
@@
Appendix B Proof of Lemma 3.3
@@
$\mathbb{E}_{k}\left[\left\lVert{\operatorname*{grad}f_{B_{k}}(x_{k})-\operatorname*{grad}f(x_{k})}\right\rVert_{2}\right]$ $=\mathbb{E}_{k}\left[\left\lVert{\frac{1}{b}\sum_{i=1}^{b}\operatorname*{grad}f_{s_{k,i}}(x_{k})-\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]$
$=\frac{1}{b^{2}}\mathbb{E}_{k}\left[\sum_{i=1}^{b}\left\lVert{\operatorname*{grad}f_{s_{k,i}}(x_{k})-\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]$
$\leq\frac{\sigma^{2}}{b},$
where the second equality comes from Assumption 3.1 (A2). From Lemma 3.2, the second term on the right-hand side of (11) yields
@@
$\mathbb{E}_{k}\left[\left\lVert{\operatorname*{grad}f_{B_{k}}(x_{k})}\right\rVert^{2}_{2}\right]\leq\frac{\sigma^{2}}{b}+\left\lVert{\operatorname*{grad}f(x_{k})}\right\rVert^{2}_{2},$
@@
Appendix D Proof of Lemma 3.6
@@
for all $k\geq 1$. From the linearity and symmetry of $P_{x_{k}}$, we obtain
@@
$\left\lVert{-\alpha_{k}P_{x_{k}}(H_{k}^{-1}g_{k})}\right\rVert^{2}_{2}\leq\alpha_{k}^{2}\left\lVert{H_{k}^{-1}g_{k}}\right\rVert_{2}\leq\alpha_{k}^{2}\nu^{2}\left\lVert{g_{k}}\right\rVert^{2}_{2},$
where the second inequality comes from $H_{k}^{-1}\preceq\nu I_{d}$. On the other hand, this inequality clearly holds if $P_{x_{k}}(H_{k}^{-1}g_{k})=0\in\mathbb{R}^{d}$. Therefore, we obtain
@@
Appendix E Linear algebra lemma
Lemma E.1.

### v2

@@ 4 Numerical experiments
We experimented with both constant and diminishing step sizes. For each algorithm, we searched in the set $\{10^{-1},10^{-2},\ldots,10^{-8}\}$ for the best initial step size $\alpha$ (both constant and diminishing). Note that the constant (resp. diminishing) step size was determined to be $\alpha_{k}=\alpha$ (resp. $\alpha_{k}=\alpha/\sqrt{k}$) for all $k\geq 1$. For each experiment, each data set was split into a training set and a test set. The experiments used a MacBook Air (M1, 2020) and macOS Monterey version 12.2 operating system. The algorithms were written in Python 3.12.1 with the NumPy 1.26.0 package and the Matplotlib 3.9.1 package. The Python implementations of the methods used in the numerical experiments are available at https://github.com/iiduka-researches/202408-adaptive.
@@ 4.1 Principal component analysis
on the Stiefel manifold $\operatorname*{St}(p,n)$. Therefore, the PCA problem can be considered to be optimization problem on the Stiefel manifold. In the experiments, we set $p$ to $10$ and the batch size $b$ to $2^{10}$. We used the QR-based retraction on the Stiefel manifold $\operatorname*{St}(p,n)$ (Absil et al., 2008, Example 4.1.3), which is defined by
@@
We evaluated the algorithms on training images of the MNIST dataset (LeCun et al., 1998) and the COIL100 dataset (Nene et al., 1996). The MNIST dataset contains $28\times 28$ gray-scale images of handwritten digits. It has a training set of 60,000 images and a test set of 10,000 images. We transformed every image into a 784-dimensional vector and normalized its pixel values to lie in the range of $[0,1]$. Thus, we set $N=60000$ and $n=784$. The COIL100 dataset contains 7,200 normalized color camera images of the 100 objects taken from different angles. As in the previous study (Kasai et al., 2019), we resized them to 32$\times$32 pixels. Moreover, we split them into an $80\%$ training set and a $20\%$ test set. Thus, we set $N=5760$ and $n=1024$.
Figure 1(a) (resp. Figure 1(b)) shows the performances of the algorithms with a constant (resp. diminishing) step size for the objective function values defined by (10) with respect to the number of iterations on the training set of the MNIST dataset, while Figures 2(a) and 2(b) present those on the test set. Moreover, Figures 5(a) and 5(b) show those on the training set of the COIL100 dataset, while Figures 6(a) and 6(b) present those on the test set. Figure 3(a) (resp. 3(b)) presents the performances of the algorithms with a constant (resp. diminishing) step size for the norm of the gradient of objective function defined by (10) with respect to the number of iterations on the training set of the MNIST dataset, while Figures 4(a) and 4(b) present those on the test set of the MNIST dataset. Moreover, Figures 7(a) and 7(b) show those on the training set of the COIL100 dataset, Figures 8(a) and 8(b) present those on the test set. The experiments were performed for three random initial points, and the thick line plots the average of all experiments. The area bounded by the maximum and minimum values is painted the same color as the corresponding line. Note that the upper part of Figures 3(a), 3(b), 4(a) and 4(b) are cut off. The initial values of the gradient norm in these figures are approximately 38, highlighting that the gradient norm was minimized during the optimization process. Furthermore, our results differ from those presented in Kasai et al. (2019) because of three key factors. In Kasai et al. (2019), the mini-batch size was fixed at 10, while in our experiments we used $2^{10}$. Furthermore, the scope of the grid search differs from that of Kasai et al. (2019). Moreover, we experimented with a doubly increasing batch size every 100 steps from an initial batch size of $b_{1}=2^{7}$ and constant and diminishing step sizes. The results are given in Appendix G.
@@
Figure 1: Objective function value defined by (10) versus number of iterations on the training set of the MNIST datasets.
@@
Figure 2: Objective function value defined by (10) versus number of iterations on the test set of the MNIST datasets.
@@
Figure 3: Norm of the gradient of objective function defined by (10) versus number of iterations on the training set of the MNIST datasets.
@@
Figure 4: Norm of the gradient of objective function defined by (10) versus number of iterations on the test set of the MNIST datasets.
(a) constant learning rate
(b) diminishing learning rate
Figure 5: Objective function value defined by (10) versus number of iterations on the training set of the COIL100 datasets.
(a) constant learning rate
(b) diminishing learning rate
Figure 6: Objective function value defined by (10) versus number of iterations on the test set of the COIL100 datasets.
(a) constant learning rate
(b) diminishing learning rate
Figure 7: Norm of the gradient of objective function defined by (10) versus number of iterations on the training set of the COIL100 datasets.
(a) constant learning rate
(b) diminishing learning rate
Figure 8: Norm of the gradient of objective function defined by (10) versus number of iterations on the test set of the COIL100 datasets.
Figures 1(a) and 2(a) indicate that RAdam and RAMSGrad (Algorithm 2) performed comparably to RASA-LR in the sense of minimizing the objective function value. Figures 1(b) and 2(b) indicate that RAdam and RAMSGrad (Algorithm 2) outperformed RASA-L and RASA-R. Figures 3(a) and 4(a) show that RAMSGrad (Algorithm 2) performed better than RASA-LR in the sense of minimizing the full gradient norm of the objective function. Figures 5(a) and 6(a) indicate that RAdam and RAMSGrad (Algorithm 2) had the best performance in the sense of minimizing the objective function value. Figures 5(b) and 6(b) indicate that RAdam and RAMSGrad (Algorithm 2) performed comparably to RASA-LR. Figures 7(a) and 8(a) shows that RAdam had the best performance in the sense of minimizing the full gradient norm of the objective function. Figures 7(b) and 8(b) indicate that RAdam and RAMSGrad (Algorithm 2) performed comparably to RASA-R and RASA-LR.
Table 1 compares performances across different batch sizes, showing the number of iterations and CPU time (in seconds) required by each algorithm with a constant step size to reduce the gradient norm below 2 when solving the PCA problem on the MNIST dataset. As our analysis (Theorem 3.5) indicates, increasing the batch size reduces the number of iterations needed to decrease the norm.
Table 1: Iterations and CPU time (seconds) required by each algorithm with a constant step size to reduce the gradient norm below 2 for solving the PCA problem on the MNIST dataset across different batch sizes. “-” indicates cases where the algorithm did not reach the threshold within the maximum allowed (1,000) iterations.
batch size $b$ number of iterations CPU time (seconds)
RSGD 256 - -
512 303 0.531
1024 149 0.394
RAdam 256 391 0.532
512 150 0.270
1024 126 0.349
RAMSGrad 256 190 0.257
512 140 0.251
1024 114 0.323
RASA-L 256 493 1.228
512 442 1.291
1024 373 1.476
RASA-R 256 397 0.662
512 371 0.809
1024 300 0.901
RASA-LR 256 - -
512 289 0.862
1024 113 0.447
Table 2 compares performances across different batch sizes, showing the number of iterations and CPU time (in seconds) required by each algorithm with a diminishing step size to reduce the gradient norm below 2 when solving the PCA problem on the MNIST dataset. As our analysis (Theorem 3.6) indicates, increasing the batch size reduces the number of iterations needed to decrease the norm. Similar tables (Tables 1116) for the remaining datasets and experiments can be found in Appendix E.
Table 2: Iterations and CPU time (seconds) required by each algorithm with a diminishing step size to reduce the gradient norm below 2 for solving the PCA problem on the MNIST dataset across different batch sizes.
batch size $b$ number of iterations CPU time (seconds)
RSGD 256 239 0.315
512 140 0.327
1024 85 0.229
RAdam 256 292 0.434
512 189 0.349
1024 101 0.280
RAMSGrad 256 234 0.324
512 224 0.437
1024 141 0.384
RASA-L 256 409 1.008
512 370 1.133
1024 282 1.102
RASA-R 256 235 0.399
512 194 0.418
1024 156 0.472
RASA-LR 256 92 0.229
512 63 0.192
1024 36 0.145
@@ 4.2 Low-rank matrix completion
We evaluated the algorithms on the MovieLens-1M [^1] https://grouplens.org/datasets/movielens/ datasets (Harper & Konstan, 2015) and the Jester [^2] https://grouplens.org/datasets/jester datasets for recommender systems. The MovieLens-1M datasets contains 1,000,209 ratings given by 6,040 users on 3,952 movies. Moreover, we split them into an $80\%$ training set and a $20\%$ test set. Thus, we set $N=4832$ and $n=3952$. The Jester dataset contains ratings of 100 jokes given by 24,983 users with scores from $-10$ to 10. In addition, we split them into an $80\%$ training set and a $20\%$ test set. Thus, we set $N=19986$ and $n=100$.
@@
Figure 9(a) (resp. Figure 9(b)) shows the performances of the algorithms with a constant (resp. diminishing) step size for objective function values defined by (12) with respect to the number of iterations on the training set of the MovieLens-1M dataset, while Figures 10(a) and 10(b) present those on the Jester dataset. Moreover, Figures 13(a) and 13(b) show those on the training set of the Jester dataset, while Figures 14(a) and 14(b) present those on the test set of the Jester dataset. Figure 11(a) (resp. 11(b)) shows the performances of the algorithms with a constant (resp. diminishing) step size for the norm of the gradient of the objective function defined by (10) with respect to the number of iterations on the training set of the MovieLens-1M dataset, while Figures 12(a) and 12(b) present those on the test set of the MovieLens-1M dataset. Moreover, Figures 15(a) and 15(b) show those on the training set of the Jester dataset, while Figures 16(a) and 16(b) present those on the test set of the Jester dataset. The experiments were performed for three random initial points, and the thick line plots the average results of all experiments. The area bounded by the maximum and minimum values is painted the same color as the corresponding line. Moreover, we experimented with a doubly increasing batch size every 20 steps from an initial batch size of $b_{1}=2^{6}$ and constant and diminishing step sizes. The results are given in Appendix G.
@@
Figure 9: Objective function value defined by (12) versus number of iterations on the training set of the MovieLens-1M datasets.
@@
Figure 10: Objective function value defined by (12) versus number of iterations on the test set of the MovieLens-1M datasets.
@@
Figure 11: Norm of the gradient of objective function defined by (12) versus number of iterations on the training set of the MovieLens-1M datasets.
@@
Figure 12: Norm of the gradient of objective function defined by (12) versus number of iterations on the test set of the MovieLens-1M datasets.
(a) constant learning rate
(b) diminishing learning rate
Figure 13: Objective function value defined by (12) versus number of iterations on the training set of the Jester datasets.
(a) constant learning rate
(b) diminishing learning rate
Figure 14: Objective function value defined by (12) versus number of iterations on the test set of the Jester datasets.
(a) constant learning rate
(b) diminishing learning rate
Figure 15: Norm of the gradient of objective function defined by (12) versus number of iterations on the training set of the Jester datasets.
(a) constant learning rate
(b) diminishing learning rate
Figure 16: Norm of the gradient of objective function defined by (12) versus number of iterations on the test set of the Jester datasets.
Figures 9(a) and 10(a) indicate that RAMSGrad (Algorithm 2) performed better than RASA-L and RASA-LR in the sense of minimizing the objective function value. Figures 9(b) and 10(b) show that RAdam and RAMSGRad (Algorithm 2) performed comparably to RASA-L and RASA-LR. Figures 11(a) and 12(a) indicate that RAdam performed comparably to RASA-R in the sense of minimizing the full gradient norm of the objective function. Figures 11(b) and 12(b) show that RAdam outperformed RASA-R. Figures 13(a), 13(b), 14(a) and 14(b) indicate that RAMSGrad (Algorithm 2) performed better than RASA-LR in the sense of minimizing the objective function value. Figures 15(a) and 16(a) indicate that RAdam performed comparably to RASA-L and RASA-R in the sense of minimizing the full gradient norm of the objective function. Figures 15(b) and 16(b) show that RAMSGrad (Algorithm 2) outperformed RASA-L and RASA-R.
@@ 5 Conclusion
This paper proposed a general framework of Riemannian adaptive optimization methods, which encapsulates several stochastic optimization algorithms on Riemannian manifolds. The framework incorporates the mini-batch strategy often used in deep learning. We also proposed RAMSGrad (Algorithm 2) that works on embedded submanifolds in Euclidean space within our framework. In addition, we gave convergence analyses that are valid for both a constant and diminishing step size. The analyses also revealed the relationship between the convergence rate and mini-batch size. We numerically compared the RAMSGrad (Algorithm 2) with the existing algorithms by applying them to principal component analysis and low-rank matrix completion problems, which can be considered to be Riemannian optimization problems. Numerical experiments showed that the proposed method performs well against PCA. RAdam and RAMSGrad performed well for constant and diminishing step sizes especially on the COIL100 dataset.
Acknowledgments
We are sincerely grateful to the Action Editor, Stephen Becker, and the three anonymous reviewers for helping us improve the original manuscript. This work was supported by a JSPS KAKENHI Grant, number JP23KJ2003.
Appendix A Useful Lemmas
Lemma A.1.
for all $k\geq 1$.
@@ Appendix A Useful Lemmas
$\mathbb{E}_{k}\left[\operatorname*{grad}f_{B_{k}}(x_{k})\right]=\frac{1}{b_{k}}\sum_{i=1}^{b_{k}}\mathbb{E}_{k}\left[\operatorname*{grad}f_{s_{k,i}}(x_{k})\right]=\operatorname*{grad}f(x_{k}).$
@@
Lemma A.2.
$\mathbb{E}_{k}\left[\left\lVert{\operatorname*{grad}f_{B_{k}}(x_{k})}\right\rVert^{2}_{2}\right]\leq\frac{\sigma^{2}}{b_{k}}+\left\lVert{\operatorname*{grad}f(x_{k})}\right\rVert^{2}_{2}$
for all $k\geq 1$.
@@
$\mathbb{E}_{k}\left[\left\lVert{\operatorname*{grad}f_{B_{k}}(x_{k})-\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]$ $=\mathbb{E}_{k}\left[\left\lVert{\frac{1}{b_{k}}\sum_{i=1}^{b_{k}}\operatorname*{grad}f_{s_{k,i}}(x_{k})-\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]$
$=\frac{1}{b_{k}^{2}}\mathbb{E}_{k}\left[\sum_{i=1}^{b_{k}}\left\lVert{\operatorname*{grad}f_{s_{k,i}}(x_{k})-\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]$
$\leq\frac{\sigma^{2}}{b_{k}},$
where the second equality comes from Assumption 3.1 (A2). From Lemma A.1, the second term on the right-hand side of (13) yields
@@
$\mathbb{E}_{k}\left[\left\lVert{\operatorname*{grad}f_{B_{k}}(x_{k})}\right\rVert^{2}_{2}\right]\leq\frac{\sigma^{2}}{b_{k}}+\left\lVert{\operatorname*{grad}f(x_{k})}\right\rVert^{2}_{2},$
@@
Appendix B Proof of Lemma 3.3
@@
for all $k\geq 1$. Here, we note that the tangent space $T_{x_{k}}M$ of an embedded submanifold $M$ in Euclidean space is a subspace of Euclidean space. Therefore, according to the result in (Axler, 2024, 6.57 (a)), the projection $P_{x_{k}}$ is a linear map. From the linearity and symmetry of $P_{x_{k}}$, we obtain
@@
$\left\lVert{-\alpha_{k}P_{x_{k}}(H_{k}^{-1}g_{k})}\right\rVert^{2}_{2}\leq\alpha_{k}^{2}\left\lVert{H_{k}^{-1}g_{k}}\right\rVert_{2}^{2}\leq\alpha_{k}^{2}\nu^{2}\left\lVert{g_{k}}\right\rVert^{2}_{2},$
where the first inequality comes from $\left\lVert{P_{x_{k}}(H_{k}^{-1}g_{k})}\right\rVert_{2}\leq\left\lVert{H_{k}^{-1}g_{k}}\right\rVert_{2}$ (see Bauschke & Combettes (2011, Corollary 3.24 (ii)) and Axler (2024, 6.57 (h)) for details) and the second inequality comes from $O\prec H_{k}^{-1}\preceq\nu I_{d}$. On the other hand, this inequality clearly holds if $P_{x_{k}}(H_{k}^{-1}g_{k})=0\in\mathbb{R}^{d}$. Therefore, we obtain
@@
Appendix C Linear algebra lemma
Lemma C.1.